OpenStudy (anonymous):

write as product: cos(6x)+ cos2(x)

OpenStudy (anonymous):

By the sum-to-product formula: cos(A) + cos(B) = 2cos[(A + B)/2]cos[(A - B)/2] With A = 6x and B = 2x: cos(6x) + cos(2x) = 2cos[(6x + 2x)/2]cos[(6x - 2x)/2] ==> cos(6x) + cos(2x) = 2cos(4x)cos(2x).

OpenStudy (cwrw238):

cos A + cos B = 2 cos (A+B) / 2 * sos (A - B) / 2

OpenStudy (anonymous):

\[\cos(C) + \cos(D) = 2 \cos(\frac{C + D}{2}) \cdot \cos(\frac{C - D}{2})\]

OpenStudy (anonymous):

ohh, i forgot about that formula, thanks :)